If we were to add a piece of solid Cu to an aqueous solution of silver (1) nitrate, the Silver would be replaced in a single replacement reaction forming aqueous copper (11) nitrate and solid silver. How much silver is produced if 15.00 grams of Cu is added to the solution of ecxess silver nitrate.


1
Expert's answer
2021-03-23T05:54:29-0400

Solution:

The balanced chemical equation:

Cu(s) + 2AgNO3(aq) → Cu(NO3)2(aq) + 2Ag(s)

According to the equation above: n(Cu) = n(Ag)/2


Moles of Cu = Mass of Cu / Molar mass of Cu

The molar mass of Cu is 63.546 g mol-1.

Hence,

Moles of Cu = 15.00 g / 63.546 g mol-1 = 0.236 mol

n(Cu) = 0.236 mol


n(Ag) = 2 × n(Cu) = 2 × 0.236 mol = 0.472 mol

Mass of Ag = Moles of Ag × Molar mass of Ag

The molar mass of Ag is 107.868 g mol-1.

Hence,

Mass of Ag = 0.472 mol × 107.868 g mol-1 = 50.9137 g = 50.9 g

Mass of Ag = 50.9 g


OR (short form solution):

(15.00 g Cu) × (1 mol Cu / 63.546 g Cu) × (2 mol Ag/1 mol Cu) × (107.868 g Ag / 1 mol Ag) = 50.9 g Ag


Answer: 50.9 grams of silver (Ag) is produced.